OpenStudy (musicequalslife):

what is the missing term in the arithmetic sequence 6, __,30

OpenStudy (anonymous):

this being an arithmetic sequence, we need to find what is called arithmetic mean .. Do you know what that is ?

OpenStudy (musicequalslife):

No not really aha im not really good with math but I think its 18 its multiple choice

OpenStudy (musicequalslife):

12, 18, 22, 15

OpenStudy (anonymous):

lets solve this problem let the sequence be a,a+d,a+2d where a is the first term and d is the common difference does this make sense ?

OpenStudy (musicequalslife):

umm no lol i'm so sorry do you mean like 6+30 and 6+2(30)?

hartnn (hartnn):

in arithmetic sequence, the difference between next term and current term is same let your middle term be x so x-6 = 30-x get x from here

OpenStudy (dayakar):

1,3,5,7,9,... here each term comes when add a fixed number except first one means a1 =1 a2= 1+2 =3 (second term) a3= 3+2 =5

OpenStudy (musicequalslife):

so it would be 18 I did what you said @hartnn and I got 18

OpenStudy (anonymous):

I mean the sequence be a, a+d, a+2d... which means a= 6 a+2d= 30 using this you can find d... what you need is a+d :) Can you solve it from here ?

hartnn (hartnn):

good! \(\huge \checkmark \)

OpenStudy (dayakar):

a,b,c are in A,P b= (a+c)/2

OpenStudy (musicequalslife):

Thank you, you guys!!! <333

OpenStudy (dayakar):

6 ,b,30 b = (6+30)/2 = 36/2= 18

OpenStudy (musicequalslife):

Thank you all for helping :)